WLCG-OSG-EGEE Ops' Minutes Mon 09 Feb 2009

Summary

  • A plan was presented for rolling out recommendations of the installed capacity document. This included new information providers for storage, DPM, dCache, CASTOR and STORM. For CEs a new yaim will just add a couple of extra values to publish in the GlueSchema.

Attendance

EGEE

  • Asia Pacific ROC: ShuTing Liao
  • Central Europe ROC: Malgorzata
  • OCC / CERN ROC: John Shade, Antonio Retico, Nick Thackray, Steve Traylen
  • French ROC:
  • German/Swiss ROC: Wen Mei
  • Italian ROC: Paolo
  • Northern Europe ROC: Jens Larsson, Gert Svensson
  • Russian ROC: Victor, Lev
  • South East Europe ROC: Kostas
  • South West Europe ROC: Kai
  • UK/Ireland ROC: Jeremy
  • GGUS: Maria, Torsten
  • GOCDB:

WLCG

  • WLCG Service Coordination: Harry Renshall
  • ASGC: Jason
  • BNL: Absent
  • CERN site: Ulrich
  • FNAL: Catalin
  • FZK: Angela
  • IN2P3:
  • INFN:
  • NDGF: Jens Larson, Roger Oscarsson, Leif
  • PIC:
  • RAL: Gareth
  • SARA/NIKHEF: Ron
  • TRIUMF:

LHC Experiments

  • ATLAS: Alessandro di Girolamo
  • LHCb: Roberto Santinelli
  • CMS: absent
  • ALICE: absent

Feedback on Last Week's Minutes

None was given.

EGEE Items

Grid Operator Hand Over on Duty

  Primary Team Secondary Team
From Italy France
To Russia UK/I

  • GGUS:45181 CERN-ROC Site, SDU. Will be suspended after the meeting if no obvious activity.
    • Site has been contacted but they will be suspended on Wednesday unless something happening.
  • UK/I one person down today for ROC duty. Helen has approved just now the extra person, UK/I should be up and running.

Sites Considered For Suspension

PPS Reports and Issues

  • OpsMeetingPps
  • Staged roll out across ROCs.
    • Services that still need to be covered , BDII covered by RAL , CE covered by Central Europe. We need volunteers to roll out services. LCG.PPS_Release_Testing
  • SA3 releases to everyone pending on this controlled rollout.
  • GLExec and SCAS Pilot - > FZK, Lyon and NIKHEF. Pilot process in April. PpsPilotSCAS
  • PpsPilotCream underway and previous failiures are progressing, 40 jobs/minute can now be sustained.

gLite Release News

  • OpsMeetingGliteReleases
  • Updates to gLite 3.0 and 3.1. Contains voms certificate updates. We realise that we forgot the WMS 3.0 from the list of affected services. But you should upgrade anyway.
  • glite 3.1 update 43 next release.
    • Update to WMS, last update before a pilot WMS service.
    • Info provider to SGE.
    • New BDII being tested at RAL.

EGEE Items From ROC Reports

$ Italy: INFN-T1 FTM endpoint (FTM endpoint at CNAF):

WLCG Items

WLCG issues coming from ROC reports

  • None.

Upcoming WLCG Service Interventions

  • Consult links on the agenda page.

Plan for rolling out the recommendations of the WLCG Installed Capacity document

WLCG_GlueSchemaUsage-1.8.pdf

  • Approved at last weeks management board.
  • Rollout at end of February.
    • Either the old SpecInt2000 or the HEPSpec2006.
  • Storage.
    • New providers for DPM are in certification.
    • New providers for dCache are being written.
    • New Storm one with v1.4.0.
  • Monitoring
    • gstat & gridmap will be updated, times will be coming.
Kostas
Is this a big upgrade to do quickly.
Flavia
It is just an info provider and/or YAIM update with a couple of new parameters.
Steve
For the CE it is not a big update particularly.
Flavia
DPM/Castor providers already deployed at some sites.

Wiki page containing FTM Endpoints

Can all tier-1 sites please keep the list of FTM endpoints up to date. The list is here: LCGFTMEndpoints Please T1s can you update this page.

Please ATLAS can you provide your list of updates.

WLCG Service Coordination

Atlas will be deleting lots of redundant files over the next couple of weeks. Rate will be slow.

ATLAS Service

  • Have checked SRMv2 tests against BNL site that.
  • We are waiting on the LFC endpoint publication for the BNL LFC.
    • Why not GGUS. There is a ticket about sorting out ticketing from GGUS <-> BNL at the moment.
  • All VOBoxes have been updated and now also being published into CERN's SLS.

ALICE Service

Not present.

CMS Service

Not present.

LHCb Service

Not present.

OSG Items

  • A reminder was sent to GGUS developer for progress on GGUS-OSG ticket update flow testing based on GGUS:45488
    • In progress.
  • Rob or Kyle or other OSG supporter to re-launch Felipe Silva on GGUS:45094.
    • The submitter does not accept the suggestion to close the ticket.

Newly Created Action Items

Review of Open Action Items

Vo Space
Reported for LHC VOs, Close ticket.
RedIris
Steve will have a look.
Libs needed for 32bit -> 64bit migration
There is no list. There is a list of per VO on the VO Cards, we may try and produce a common list.
FTM Endpoints
Wikipage exists and this action item is closed.

Open Action Items

IdSubmitterDescriptionCreationDueAssigned To 

Actions Closed in Last 20 Days

IdSubmitterDescriptionCreationDueAssigned ToClosed 

AOB

Next Meeting

The next meeting will be Monday, 16 Feb 2009 15:00 UTC (16:00 Swiss local time).

  • Attendees can join from 14:45 UTC (15:45 Swiss local time) onwards.
  • The meeting will start promptly at 15:00 UTC (16:00 Swiss local time).
  • The WLCG section will start at the fixed time of 15:30 UTC (16:30 Swiss local time).
  • To dial in to the conference:
    • Dial +41227676000
    • Enter access code 0148141


These minutes can only be changed by members of:

Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2009-02-12 - SteveTraylen
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    EGEE All webs login

This site is powered by the TWiki collaboration platform Powered by Perl This site is powered by the TWiki collaboration platformCopyright & by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Ask a support question or Send feedback